OpenStudy (anonymous):

inequalities 2/3(x+3/4)<5/6

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }(\frac{ x+3 }{ 4 })< \frac{ 5 }{ 6 }\] multiply both sides by 3/2 to get rid of 2/3:\[\frac{ x+3 }{ 4 }<\frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

can you do the rest yourself?

OpenStudy (anonymous):

yes but i dont know what to do with the fraction @mihirb

OpenStudy (anonymous):

which fraction?

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

@mihirb

OpenStudy (mysangels):

fractions are just division, so you divide 5 by 4

OpenStudy (mysangels):

but first get rid of x +3 with x+3-3

OpenStudy (anonymous):

no @mysangles

OpenStudy (anonymous):

i mean no @mysangels

OpenStudy (mysangels):

then divide that answer with 4, and get the answer. just remember that the sign doesn't have to do anything with solving the problem

OpenStudy (anonymous):

multiply both sides by 4

OpenStudy (anonymous):

then u get:\[x+3<5\]

OpenStudy (anonymous):

and now you can solve it

OpenStudy (mysangels):

oh, yes you are right. that is how you solve the inequality

OpenStudy (anonymous):

the answer is x<2 right @mihirb